OpenStudy (anonymous):

Find the zeros of the function: f(x) = x2 - 5x - 6. Enter the smallest zero value. Use a number only.

OpenStudy (anonymous):

X=6 or X= -1

Directrix (directrix):

x^2 - 5x - 6 = 0 (x - 6) (x + 1) = 0 x = 6 or x = -1 Enter the smallest zero value. ---> -1